# Exploring Power Side-Channel Challenges in Embedded Systems Security

### v1

@@ Abstract
Power side-channel (PSC) attacks are widely used in embedded microcontrollers, particularly in cryptographic applications to extract sensitive information. However, expanding the applications of PSC attacks to broader security contexts in the embedded systems domain faces significant challenges. These include the need for specialized hardware setups to manage high noise levels in real-world targets and assumptions regarding the attacker’s knowledge and capabilities. This paper systematically analyzes these challenges and introduces a novel signal-processing method that addresses key limitations, enabling effective PSC attacks in real-world embedded systems without requiring hardware modifications. We validate the proposed approach through experiments on real-world black-box embedded devices, verifying its potential to expand its usage in various embedded systems security applications beyond traditional cryptographic applications.
@@ 1 Introduction
Power side-channel (PSC) attacks are typically used to extract secrets from embedded Microcontroller Units (MCUs). These attacks [1, 2, 3, 4] exploit information leakage from power usage of the embedded devices. The leakage is primarily due to the physical characteristics of electronic components, exhibiting varying behaviors depending on the logical values they process. Although side-channel attacks in embedded systems were initially developed to target encryption algorithms [5], in recent years, researchers have broadened the application of the PSC to other embedded systems security domains, including instruction disassembly [6, 7, 8], code execution tracking [9], fuzzing [10], and intrusion detection [11].
Despite significant advances, various challenges still limit the applications of PSC in embedded system security. Notably, PSC attacks on cryptographic algorithms often assume a well-equipped attacker with a specialized target board designed for PSC measurements. Furthermore, it is generally presumed that the attacker is aware of the cryptographic algorithm’s implementation and has access to the target device or a similar one to study it before launching the attack. However, these assumptions do not always hold in other contexts, such as instruction disassembly, where the attacker may not know the executed algorithm. Instruction disassembly typically relies on thousands of measurements (templates) for later analysis. Regarding instruction disassembly, existing research has focused on MCUs with limited bus width (e.g., 8-bit or 16-bit MCUs) or simple CPU pipelines. We argue that these MCUs no longer represent realistic targets for state-of-the-art embedded systems security research.
Similarly, research on control-flow reconstruction and execution tracking is often evaluated on the platform using limited-bus-width MCUs or evaluation environments that have no peripheral interactions. This is not a realistic assumption but is evaluated in this way to reduce the noise. The same pattern is observed in fuzzing applications [10], where researchers gather power traces in an offline setup and employ machine learning algorithms for post-processing at a later time. This method does not align with real-time fuzzing tasks, as the data is not processed or transmitted in real-time to the target application.
In the intrusion detection research [11], evaluation platforms often feature hardware modifications to reduce noise. For instance, in the existing literature [12], the evaluation target for an Ardupilot controller does not have any peripherals, limiting the practicality of using PSC for flight control software which is designed to interact with their peripherals to control the drone.
Based on the above observations, we argue that most research in extending the application of PSC beyond traditional cryptographic applications (e.g., [10, 11, 8]) contains unrealistic assumptions and methodologies. Therefore, we do not intend to investigate evaluations for their proposed solutions, as done in other fields [13]. Instead, this work aims to identify key challenges in PSC applications that lead to such practices and suggest novel solutions. These challenges may be general, such as the presence of noise, or they might be specific to certain applications (e.g., real-time requirements in fuzzing).
In this paper, we explore the application of PSCs in embedded systems by first discussing potential uses in software security and identifying the associated challenges. We systematically analyze existing PSC-based threat models within the embedded systems domain and explore why current solutions struggle to generalize across different contexts. This analysis includes assessing the effectiveness of these solutions on real-world black-box embedded targets. To address these challenges, we introduce a novel method based on signal processing algorithms. Notably, our method is the first to work on real-world black-box embedded devices without requiring any hardware modifications (e.g., shunt resistor or high-resolution ADC) for side-channel probes. Additionally, our method generalizes across various board designs without the need for target profiling. This work allows further usage of PSC in embedded system security, enabling previously impossible applications. To foster research on this topic, we release the experimental data sets and our signal processing algorithm implementation on Github [14].
In summary, we make the following contributions:
@@
  • We perform a first comprehensive and systematic analysis of the existing PSC applications for embedded devices software security.
@@
  • We identify ten key conditions of the attacker models and five applications that exploit PSC in embedded devices.
@@
  • We propose a novel non-profiling solution based on signal processing to address existing challenges and evaluate them on real-world targets.
@@ 2.2 Frequency Domain Transformations
PSC traces are time series that capture the changes in voltage over time during the execution of a process. The time domain representation of these traces shows how the signal evolves as time progresses. This view helps analyze how the signal changes at each moment and for visualizing patterns. However, analyzing signals purely in the time domain can be computationally expensive and may not reveal important characteristics of the signal. To make signal analysis more efficient and informative, we can transform the signal from the time domain to the frequency domain. This transformation, commonly done using methods like the Fourier Transform [15], breaks the signal down into its component frequencies. In other words, it shows which frequencies are present in the signal and how much each frequency contributes to the overall signal. This frequency-based representation provides a different, often more intuitive, view of the signal, allowing for better signal processing and analysis by highlighting patterns that are not as obvious in the time domain. One significant advantage of working in the frequency domain is that certain operations are simplified in this domain (for example, convolution in the time domain is transformed into a multiplication in the frequency domain). Furthermore, various signal-processing techniques are specifically designed for the frequency domain. For instance, the wavelet transform [16] is a well-known method used to represent a signal in terms of both the existing frequencies and a ”wavelet” function. In simple terms, the wavelet transform scans the time-series, looking for repeating patterns that match the base wavelet function’s frequency. This makes the wavelet transform particularly powerful because it analyzes the frequencies in a signal and retains some information about the original time series. Unlike other methods, such as the Fourier transform, which focuses purely on frequencies and loses the connection to when those frequencies occur, the wavelet transform allows you to see both the frequency content (the different frequencies that make up the signal) and where in the time series those frequencies are present.
@@ 3 General Limitations
To understand the general limitations of PSC attacks on embedded devices, we start by examining the common challenges and reviewing the solutions proposed by the research community. Some limitations are observable across various applications, we focus on evaluating these application-independent limitations and discuss how they affect the overall effectiveness of PSC attack on embedded systems.
@@ 3.1 Probing
Figure 1: The ChipWhisperer target board with an integrated shunt resistor on the PCB. This resistor enables accurate measurement of power consumption, which is the current method for conducting precise PSC analysis [17].
PSCs are usually measured using a shunt resistor (1 to 100 Ohm, depending on the current consumption of the processor) on GND (ground) or main power supply (VCC or VDD) pins. A shunt resistor is a resistor placed on the power supply lines of the target processor, causing a small voltage drop between the main power supply lines and the processor. The voltage drop caused by the processor’s power consumption, which is different for various bit sequences, is the main source of PSC leakage. The ideal probing point for PSC leakage is after the shunt resistor. However, real-world circuit boards targeted by PSC attacks usually don’t have any shunt resistor. Therefore, the attacker needs to either add a shunt resistor on the processor pins by manipulating the target board or measure the power traces on the main power supply of the board by adding a shunt resistor on it. In the first case, power traces are less noisy as the voltage drop does not have any effects from other components, while in the second one, the noise of other components on the Printed Circuit Board (PCB) affects the PSC traces. Hence, the probing point is very important as it significantly affects the information leakage. Consequently, as the probing point gets closer to the target processor by bypassing other components on the PCB, the amount of noise will decrease, and the measurements will have more information. Therefore, choosing a proper point for PSC probing directly affects the accuracy of an attack.
As a result, to facilitate accurate measurements, some hardware platforms have been presented. For instance, ChipWhisperer (CW)[18] is a board specifically designed for PSC attacks, featuring a soldered shunt resistor for precise measurements. Figure1 shows one of the CW target boards with an integrated shunt resistor (R3) on the PCB, strategically placed beside the main target processor to minimize noise during PSC measurements.
Another substantial element in probing is the probe itself. Generally, two types of probes are used in PSC measurements: passive probes and active probes. Passive probes do not have any external power source and are a good choice for accurate measurements with low capacitive loading, leading to accurate timing measurements. While, their performance decreases in high frequencies (more than 0.5 GHz). On the other hand, active probes require an external power supply as they have transistors for amplification. These probes are more accurate in high frequencies. For PSC measurements, passive probes are a better choice as the active probes add additional noise to the signal due to the existence of active components [19].
@@ 3.2 Noise
Known Avail. Unknown No GB2 Profiling
Known Avail. Known No WB Profiling
Fuzzing Known Avail. Known Yes WB Non Profiling
@@
TABLE I: This table outlines attacker models across various applications, categorized by the attacker capabilities. Each row represents a threat model and is categorized as white-box (WB), gray-box (GB), or black-box (BB). For example, a gray-box attacker fuzzing an embedded system is defined as a non-profiling attacker with only knowledge of how to interface with the device, lacking any knowledge of the ISA and the firmware.
Electrical noise is an inherent aspect of all electrical circuits and cannot be entirely eliminated, but its impact can be minimized depending on the application. In the context of PSC attacks, electrical noise is a crucial factor influencing scalability. The noise level in the PSC can vary based on the characteristics and components of the circuit board. Consequently, generalizing results to other boards with different designs or components, even if they use the same target processor, is often challenging or impossible in many PSC attacks. Therefore, PSC attacks are highly dependent on the specific design and characteristics of the target board. Furthermore, filtering noise from useful information is particularly challenging because PSC leakages often occur at noise levels, making it difficult to distinguish between the two. This means that attackers must not only analyze the impact of noise on the target device but also conduct extra measurements (depending on the complexity of the target and the resolution of the measurement setup) to mitigate the noise effect. In the PSC, noise can originate from various sources. Besides the noise from components, interactions with peripherals can also lead to abnormal power consumption, which can affect measurements in different ways depending on the application. Therefore, managing each type of noise requires tailored strategies to reduce its impact.
@@ 4 Challenges in Embedded Systems Security
The primary objective of this paper is to examine the application of PSC attacks in the context of embedded system security. Having discussed the general limitations, we will now explore how these limitations impact the system security domain and analyze their effects. To achieve this, we will begin by categorizing the existing applications of PSC attacks within embedded system security.
@@ 4.1 Attacker Model
  • Instruction Set Architecture (ISA): Whether the ISA is known or unknown to the attacker.
  • Firmware: Whether the attacker has the firmware image or not.
@@
  • The attacker does not know the firmware or ISA of the target. Only the semantics of the interface is known to the attacker.
  • The attacker does not have a profiling phase.
  • The attack is non-invasive, meaning that the attacker does not manipulate the target (does not add a shunt resistor).
  • The targets are real-world, off-the-shelf embedded devices without any custom design.
@@ 4.2 PSC Applications
We can envision various applications of PSC in embedded system security. Some of these applications are introduced in the literature, while others are suggested in this paper.
(a) Probing on processor’s pins (ACC=91%)
(b) Probing on board’s pins (ACC=53%)
  • Disassembly:
@@
    Disassembly attacks target executed instruction sequences in the firmware and attempt to recover it. These attacks have two different types: instruction-level and bit-level.
    Instruction-level attacks model the power consumption patterns of an instruction (template) and construct a ground truth for PSC patterns. Then, these templates can be used as a reference in the attack phase to detect the executed instruction [6]. In this context, ML models are very popular as they can generalize better than conventional template attack [7, 8]. On the other hand, bit-level approaches aim to extract executed bits inside the pipeline of the processor. Basically, each instruction is a series of bits, including opcode and operand(s). So, detecting bits can lead to full instruction recovery [31]. Moreover, register values are also targetable for both types. As a result, executed instructions can be detected based on their power consumption behavior. Table II presents an overview of the existing PSC-based disassemblers and compares them. This table explains the method employed, the target hardware, the dataset, the number of targeted instructions from the ISA, the cross-device evaluation, and the accuracy achieved. The major issue with existing disassemblers is that they don’t target complex processors. Most of them are target processors with two pipeline stages. Furthermore, the targeted applications do not include any Real-Time Operating System (RTOS). Hence, their effectiveness is not assessed on real-world applications as most of the real-world targets in the embedded system security domain run a RTOS.

### v2

@@ Abstract
Power side-channel (PSC) attacks are widely used against embedded microcontrollers, particularly in cryptographic applications, to extract sensitive information. However, expanding the applications of PSC attacks to a broader systems security context, especially in embedded systems, faces significant challenges. These challenges include the need for specialized hardware setups to manage high noise levels in real-world targets and unrealistic assumptions regarding the attacker’s knowledge and capabilities. This paper systematically analyzes these challenges and introduces a novel signal-processing method that addresses key limitations, enabling effective PSC attacks in real-world embedded systems without requiring hardware modifications. We validate the proposed approach through experiments on real-world black-box embedded devices, verifying its potential to expand its usage in various systems security applications beyond traditional cryptographic applications.
@@ 1 Introduction
Power side-channel (PSC) attacks are typically used to extract secrets from embedded Microcontroller Units (MCUs). These attacks [1, 2, 3, 4] exploit information leakage from power usage of the embedded devices. The leakage is primarily due to the physical characteristics of electronic components, exhibiting varying behaviors depending on the logical values they process. Although encryption algorithms were the initial target of side-channel attacks in embedded systems [5], in recent years, researchers have broadened the application of the PSC to other system security domains. Notably, PSC has been used for instruction disassembly [6, 7, 8] as a primary focus. Also, there are some limited applications in code execution tracking [9], offline fuzzing [10], and intrusion detection [11, 12].
When comparing PSC’s traditional use in hardware security with its potential in broader security applications, we can observe three primary areas of difference: methodology, assumptions, and evaluation.
Methodologically, research in hardware security often operates in the frequency domain, while existing work on system security applications mostly operates in the time domain. Operating in the frequency domain has significant advantages, as it allows for the filtering of noises from other hardware components operating at different frequencies. While the shift to using the frequency domain in hardware security was motivated by the requirement of advanced signal processing algorithms to defeat masking techniques designed to prevent PSC against cryptographic elements, such a transformation did not happen in system security applications, and as a result, it did not benefit from advancements in signal processing. Instead, the research in this system security often relies on time domain and pattern matching [12] to find sequences of signals in different time windows, which is inefficient.
Additionally, the assumptions that are being made between hardware security and system security are different. In the context of hardware security, it is generally presumed that the attacker is aware of the cryptographic algorithm’s implementation and has access to the target device or a similar one to study it before launching the attack. While these prerequisites might be acceptable in certain system security applications (such as designing our PCB for intrusion detection), they may not be suitable for other applications (e.g., fuzzing a third-party board). Therefore, the profiling-based methods face challenges in generalizing across different PCBs, especially for black-box embedded systems (i.e., embedded systems in which their firmware, source code, or Instruction Set Architecture (ISA) is inaccessible).
Finally, evaluation techniques in hardware security-focused PSC research differ significantly from those in the system security domain. Evaluations in hardware security are often conducted under highly controlled conditions, frequently using custom-built boards designed to minimize hardware noise [13]. Such setups allow researchers to test the effectiveness of security measures like algorithmic masking techniques [14]. However, these controlled environments do not accurately reflect real-world conditions in system security, where noise and variability are substantial factors.
Interestingly, some system security applications of PSC adopt similarly controlled approaches for their evaluation. For instance, evaluations are often carried out on isolated MCUs rather than on integrated PCBs [9, 8], sometimes using low-end microcontrollers [15, 16] (e.g., 8-bit or 16-bit) or boards modified to further reduce noise, such as through the addition of shunt resistors [11]. In these setups, peripheral interactions are minimized or do not exist to limit hardware noise even further. In contrast, power traces in real-world targets are considerably noisier, incorporating interference from various hardware (peripheral interactions, hardware components on the board) and firmware components (e.g. context switches) and running on more complex MCUs.
Based on the above observations, we argue that most research in extending the application of PSC beyond traditional cryptographic applications contains unrealistic assumptions, methodologies, and evaluation. Inspired by the work of Muench et al. [17] on addressing challenges in fuzzing embedded systems firmware, our objective in this work is twofold:
First, we identify the existing challenges that lead to impractical methodologies, assumptions, and evaluations in PSC applications within the system security context [10, 11, 8]. These challenges may be broad (such as the pervasive issue of noise), or they may pertain to specific applications, like real-time requirements in fuzzing. Furthermore, we explore the application of PSCs in system security. As a result, we systematically analyze existing PSC-based attacker models within the embedded systems domain and explore why current research struggles to generalize across different contexts.
Second, we propose a novel signal-processing algorithm specially designed for system security applications to process the power traces in the frequency domain and suppress noise both in hardware and software. Notably, our method is the first to work on real-world black-box embedded devices without requiring any hardware modifications (e.g., shunt resistor or high-resolution ADC) for side-channel probing. Additionally, our method generalizes across various board designs without the need for target profiling or performing template attacks, something that does not exist today.
To the best of our knowledge, we are the first to apply power side-channel on high-end embedded systems (an application processor CPU clocked at 1Ghz with a Cortex-A 64bit ARM CPU instead of traditional Microcontrollers). As a result of this novel algorithm, we explore additional improvements to existing system security applications, namely fuzzing and intrusion detection, and introduce another application, namely software composition analysis, which allows us to detect algorithms running inside real-world embedded devices without access to the firmware or performing template attack to model the MCU. We will conduct thorough evaluations to demonstrate these results and underscore the potential for further extending the application of PSC across additional system security domains.
To foster research on this topic, we release the experimental data sets and our signal processing algorithm implementation on Github [18]. In summary, we make the following contributions:
@@
  • We perform a first comprehensive and systematic analysis of the existing PSC applications for embedded devices system security.
@@
  • We identify ten key attacker models and five applications that exploit PSC in embedded devices.
@@
  • We propose a novel solution based on signal processing to address existing challenges and evaluate them on real-world targets, including high-end processors.
@@ 2 Background
(a) Raw Power Trace in Time Domain
(b) FFT and Bandpass Filter
(c) Filtered Power Trace
Figure 1: Frequency domain representation shows a better view of the signal that can help distinguish power leakage originating from different sources of the target board.
@@ 2.2 Frequency Domain Transformations
PSC traces are time series that capture the changes in voltage over time during the execution of a process. The time domain representation of these traces shows how the signal evolves as time progresses. This view helps better pattern visualization and signal analysis. However, analyzing signals purely in the time domain can be computationally expensive and may not reveal important characteristics of the signal. To make signal analysis more efficient and informative, we can transform the signal from the time domain to the frequency domain. This transformation is commonly done using methods like the Fourier Transform [19], which breaks the signal down into its component frequencies. In other words, it shows which frequencies are present in the signal and how much each frequency contributes to the overall signal. This frequency-based representation provides a different, often more intuitive, view of the signal, allowing for better signal processing and analysis by highlighting patterns that are not as obvious in the time domain. One significant advantage of working in the frequency domain is that certain operations are simplified in this domain (for example, convolution in the time domain is transformed into a multiplication in the frequency domain). As an example, we recorded the power trace of a Raspberry Pi 4B while writing to memory. As we show in Figure 1 (a), the power trace has massive peaks (marked with red stars) that disturb the signal, while the smaller peaks in between are due to memory interaction. However, by transforming the signal to the frequency domain (Figure 1 (b)), we distinguish those peaks and thus filter them using a bandpass filter. Then, we return the filtered signal to the time domain (Figure 1 (c)), which shows a clean signal with only memory interactions. Furthermore, various signal-processing techniques are specifically designed for the frequency domain. For instance, the wavelet transform [20] is a well-known method used to represent a signal in terms of both the existing frequencies and a ”wavelet” function. In simple terms, the wavelet transform scans the time series, looking for repeating patterns that match the base wavelet function. This makes the wavelet transform particularly powerful because it analyzes the frequencies in a signal and retains some information about the original time series. Unlike other methods, such as the Fourier transform, which focuses purely on frequencies and loses the connection to when those frequencies occur, the wavelet transform allows you to see both the frequency content (the different frequencies that make up the signal) and wherein the time series those frequencies are present.
2.3 Low-end and High-end Embedded Devices
Embedded devices are divided into two categories: low-end and high-end. Low-end devices typically have a microcontroller unit (MCU) and a single core with RAM integrated inside the chip. On the other hand, high-end devices refer to those with a microprocessor unit (MPU). MPUs usually have multiple complex cores with cache and external RAM. They operate at higher frequencies compared to MCUs and have more complex architectures. Existing work in PSC for system security focuses on low-end targets, while nowadays high-end embedded devices are widespread.
@@ 3 General Limitations
To understand the general limitations of PSC attacks on embedded devices, we start by examining the common challenges and reviewing the solutions proposed by the research community. Some limitations are observable across various applications. We evaluate these application-independent limitations and discuss how they affect the effectiveness of PSC attacks on embedded systems.
@@ 3.1 Probing
PSCs are usually measured using a shunt resistor (1 to 100 Ohm, depending on the current consumption of the processor) on GND (ground) or main power supply (VCC or VDD) pins. A shunt resistor is a resistor placed on the power supply lines of the target processor, causing a small voltage drop between the main power supply lines and the processor. The voltage drop caused by the processor’s power consumption, which is different for various bit sequences, is the main source of PSC leakage. The ideal probing point for PSC leakage is after the shunt resistor. However, real-world circuit boards targeted by PSC attacks usually don’t have any shunt resistor. Therefore, the attacker needs to either add a shunt resistor on the processor pins by manipulating the target board or measure the power traces on the main power supply of the board by adding a shunt resistor on it. In the first case, power traces are less noisy as the voltage drop does not have any effects from other components, while in the second one, the noise of other components on the Printed Circuit Board (PCB) affects the PSC traces. Hence, the probing point is very important as it significantly affects the information leakage. Consequently, as the probing point gets closer to the target processor by bypassing other components on the PCB, the amount of noise will decrease, and the measurements will have more information. Therefore, choosing a proper point for PSC probing directly affects the accuracy of an attack. As a result, to facilitate accurate measurements, some hardware platforms have been presented. For instance, ChipWhisperer (CW) [21] is a board specifically designed for PSC attacks, featuring a soldered shunt resistor for precise measurements.
@@ 3.2 Noise
Known Available Unknown No GB2 Profiling
Known Available Known No WB Profiling
Fuzzing Known Available Known Yes WB Non Profiling
@@
TABLE I: This table outlines attacker models across applications, categorized by the attacker capabilities. Each row represents a threat model and is categorized as white-box (WB), gray-box (GB), or black-box (BB). For example, a gray-box attacker fuzzing an embedded system is defined as a non-profiling attacker with only knowledge of how to interface with the device, lacking any knowledge of the ISA and the firmware.
Electrical noise is an inherent aspect of all electrical circuits and cannot be eliminated, but its impact can be minimized depending on the application. In the context of PSC attacks, electrical noise is a crucial factor influencing scalability. The noise level in the PSC can vary based on the characteristics and components of the circuit board. Consequently, generalizing results to other boards with different designs or components, even if they use the same target processor, is often challenging or impossible in many PSC attacks. Therefore, PSC attacks are highly dependent on the specific design and characteristics of the target board. Furthermore, filtering noise from useful information is particularly challenging because PSC leakages often occur at noise levels, making it difficult to distinguish between the two. This means that attackers must not only analyze the impact of noise on the target device but also conduct extra measurements (depending on the complexity of the target and the resolution of the measurement setup) to mitigate the noise effect. In the PSC, noise can originate from various sources. Besides the noise from components, interactions with peripherals can also lead to abnormal power consumption, which can affect measurements in different ways depending on the application. Therefore, managing each type of noise requires tailored strategies to reduce its impact.
@@ 4 Challenges in Embedded Systems Security
The primary objective of this paper is to examine the application of PSC attacks in the context of embedded systems security. Having discussed the general limitations, we will now explore how these limitations impact the system security domain and analyze their effects (Addressing RQ2). To achieve this, we will begin by categorizing the existing applications of PSC attacks within embedded system security.
@@ 4.1 Attacker Model
  • ISA: Whether the ISA is known or unknown to the attacker.
  • Firmware: Whether the firmware image is available to the attacker or not.
@@
  • The attacker does not know the firmware or ISA of the target, he knows only the semantics of the interface.
  • The attacker does not have to profile the same device.
  • The attack is non-invasive, meaning that the attacker does not manipulate the target (no shunt resistor).
  • The targets are real-world, off-the-shelf embedded devices without any custom design for PSC measurement.
@@ 4.2 PSC Applications
We can envision various applications of PSC in embedded system security. Some of these applications are introduced in the literature, while others are suggested in this paper. This section addresses RQ1.
4.2.1 Disassembly:
@@
Disassembly attacks focus on executed instruction sequences within firmware, aiming to recover its contents. These attacks can be classified into two types: instruction-level and bit-level
Instruction-level attacks involve modeling the power consumption pattern of individual instructions to create a reference template for PSC patterns. During the attack phase, the attacker leverages these templates to identify the executed instructions [6]. ML models outperform other methods in this context, as they often generalize more effectively than traditional template-based approaches [7, 8]. bit-level approaches, by contrast, focus on extracting bits of state from the processor pipeline. Each instruction consists of a sequence of bits that includes the opcode and operand(s); detecting these bits can potentially reveal the entire instruction [31]. Table II provides an overview of existing PSC-based disassemblers, comparing their methodologies, target hardware, datasets, the number of instructions targeted from the ISA, cross-device evaluation, and accuracy. A key limitation of current disassemblers is their focus on simple processors, usually with just two pipeline stages, and their lack of testing on applications with a Real-Time Operating System (RTOS), which is common in real-world embedded systems. Consequently, their effectiveness in real-world scenarios remains largely unassessed. As outlined in Table I, the attacker model for these studies generally assumes a gray-box approach, where the attacker has knowledge of the ISA and firmware during profiling. However, this model does not extend well to black-box scenarios, where such information is unknown to the attacker.
4.2.2 Control Flow Graph: